The difference of two numbers is 14. three times the larger number is 45 less than four times the smaller. What are the two numbers.
Answer by mananth(16946)

x y the numbers x>y
x-y=14...............1
3x=4y-45
3x-4y=45..............2
multiply (1) by -1 and add the equation to (2)
-3x+3y=-42
3x-4y=45
-y=3
y=-3
..
plug y in (1)
x-(-3)=14
x+3=14
x=11